The coin size data (measured in millimeters) collected from each group is shown below.

Low Income   High Income
28   19
17   18
26   21
18   17
24   21
16   25
14   11
33   24
13   16
24   21
26   18
31   30
23   20
27   20
18   21
25  
14  
16  
27  
31  


You can copy the data into Excel by highlighting the data, right-clicking and selecting Copy, then opening Excel, clicking on a blank cell, and selecting Paste from the Edit menu.

Give your answers to at least 3 decimal places accuracy. Keep the data in Excel, as you will need this information again in later questions.

2. Compute the following summary statistics

a) The mean for the low income group is:


b) The median for the low income group is:


c) The standard deviation for the low income group is:


d) The mean for the high income group is:


e) The median for the high income group is:


f) The standard deviation for the high income group is:
Answer #1


Answer Date: 02/10/2019 a) The mean of low income is, Xx 28 +17+.......... +17+31 20 = 22.550 The mean of low income is 22.5513 23 28 | | 14 24 31 | | 14 24 31 16 25 33 16 26 17 26 18 27 18 27 | | | | M M10 + M11 = 24+24 = 24.000 The median of low inThe standard deviation of low income is, 2(x - 7)? V n-1 (28 - 22.550)+(17-22.550) + ..... +(27-22.550) +(31 - 22.550) 20-1 =The mean of high income is 20.133 The median of high income is, First, rearrange the data from smallest to largest then findThe standard deviation of high income is, 2(x – 7) V n-1 (19-20.133)2 +(18-20.133)2 +..... +(20 - 20.133)2 +(21- 20.133)2 15-